# DORNet: A Degradation Oriented and Regularized Network for Blind Depth Super-Resolution

### v3

@@ 1 Introduction
Blind depth super-resolution (DSR) aims to recover precise high-resolution (HR) depth from low-resolution (LR) depth with unknown degradation, which has been widely applied in many fields, such as virtual reality [29, 16, 40, 32], augmented reality [30, 42, 39], and 3D reconstruction [3, 4, 41, 43]. As shown in Fig. 1(a), recent RGB-guided DSR methods [46, 52, 34, 25, 2] have been proposed that integrate RGB features aligned with input depth based on the assumption of known and fixed degradation, achieving excellent performance on synthetic data.
However, due to limitations in sensor technology and imaging environments, depth data captured from real-world scenes often suffer from unconventional and unknown degradation (e.g., structural distortion and blur). Such real-world degradation results in structure inconsistency between depth and RGB, impairing the performance of previous methods. Moreover, real-world degradation labels are unavailable, preventing us from explicitly estimating the degradation between LR depth and HR depth.
@@
To address these issues, as shown in Fig. 1(b), we propose a degradation oriented and regularized network (DORNet). The DORNet utilizes degradation representations to guide the restoration of HR depth maps from real-world scenarios with unknown degradation. To this end, we present a self-supervised degradation learning strategy to estimate the implicit degradation representations between LR and HR depth. In this strategy, a router mechanism is first introduced to dynamically control the generation of degradation kernels with varying scales. We then design degradation regularization that leverages these kernels to deteriorate the predicted HR depth, yielding a new degraded depth. Consequently, the entire degradation process is learned by narrowing the distance between the new degraded depth and the LR depth, without using degradation labels. Furthermore, we observe that the HR RGB image can provide sharp and complete details for the degradation areas of the LR depth. Therefore, we propose utilizing the estimated degradation to adaptively select RGB features to guide and facilitate the RGB-D fusion. Concretely, we develop a degradation-oriented fusion scheme, deploying a degradation-oriented feature transformation module (DOFT). The DOFT produces filter kernels from the learned degradation and then filters the RGB features, offering complementary contents for the depth features.
@@
Figure 3: Overview of DORNet. Given $\boldsymbol{D}_{up}$ as input, the degradation learning first encodes it to produce degradation representations $\boldsymbol{\tilde{D}}$ and $\boldsymbol{D}$. Then, $\boldsymbol{\tilde{D}}$, $\boldsymbol{D}$, $\boldsymbol{D}_{lr}$, and $\boldsymbol{I}_{r}$ are fed into multiple degradation-oriented feature transformation (DOFT) modules, generating the HR depth $\boldsymbol{D}_{hr}$. Finally, $\boldsymbol{D}$ and $\boldsymbol{D}_{hr}$ are sent to the degradation regularization to obtain $\boldsymbol{D}_{d}$, which is used as input for the degradation loss $\mathcal{L}_{deg}$ and the contrastive loss $\mathcal{L}_{cont}$. The degradation regularization only applies during training and adds no extra overhead in testing.
@@ 3.1 Network Architecture
Given LR depth $\boldsymbol{D}_{lr}\in R^{h\times w\times 1}$ with unknown degradation and RGB $\boldsymbol{I}_{r}\in R^{sh\times sw\times 3}$ as inputs, our method aims to recover accurate HR depth $\boldsymbol{D}_{hr}\in R^{sh\times sw\times 1}$ by learning the degradation representations. $h$, $w$, and $s$ represent the height, width, and upsampling factor, respectively.
As shown in Fig. 3, our DORNet mainly consists of a self-supervised degradation learning strategy (green part) and a degradation-oriented fusion scheme (orange part). Specifically, the upsampled LR depth $\boldsymbol{D}_{up}\in R^{sh\times sw\times 1}$ is first input into the degradation learning, producing both the router and the degradation representations, $\boldsymbol{\tilde{D}}$ and $\boldsymbol{D}$. Then, $\boldsymbol{\tilde{D}}$, $\boldsymbol{D}$, $\boldsymbol{D}_{lr}$, and $\boldsymbol{I}_{r}$ are sent to multiple degradation-oriented feature transformation modules (DOFT), which selectively propagate RGB information into the depth features, resulting in HR depth $\boldsymbol{D}_{hr}$. Next, the degradation regularization takes $\boldsymbol{D}$ as input and utilizes routing selection to adaptively generate degradation kernels with varying scales, all of which are sent into the filtering and summation modules together with $\boldsymbol{D}_{hr}$, obtaining the new degraded depth $\boldsymbol{D}_{d}$. Finally, $\boldsymbol{D}_{d}$ is employed as input for the degradation loss $\mathcal{L}_{deg}$ and the contrastive loss [36] $\mathcal{L}_{cont}$, further promoting the learning of degradation representations.
@@ 3.3 Degradation-Oriented Fusion
As shown in the orange part of Fig. 3, $\boldsymbol{D}_{lr}$ is first input into bicubic upsampling. Then, the upsampled LR depth and $\boldsymbol{I}_{r}$ are mapped to $\boldsymbol{F}_{d}^{0}$ and $\boldsymbol{F}_{r}^{0}$, respectively.
@@ 4.1 Experimental Setups
Datasets. We conduct extensive experiments on both real-world RGB-D-D [8], TOFDSR [41], and synthetic NYU-v2 [26] datasets. Specifically, for the RGB-D-D, the training set comprises $2,215$ RGB-D pairs, while the test set contains $449$ pairs. Additionally, the colorization method [13] is used to fill in the raw LR depth of the TOFDC [41], obtaining the TOFDSR that includes $10K$ RGB-D pairs for training and $560$ pairs for testing. In the real-world scenarios, the LR depth is obtained using the ToF camera of the Huawei P30 Pro. Following [11, 50, 34], the synthetic NYU-v2 consists of $1,000$ RGB-D pairs for training and $449$ pairs for testing, with the LR depth generated by bicubic downsampling from the GT depth.
@@ 4.2 Comparison with the State-of-the-Art
Furthermore, Figs. 8 and 9 present the visual results on the RGB-D-D and TOFDSR datasets. In the error maps, a brighter color means a larger error. Obviously, for severely degraded LR depth, our method succeeds in recovering accurate depth structures. For instance, the handbag in Fig. 8 predicted by our method is more precise than others. Additionally, the error maps in Fig. 9 show that DORNet reconstructs HR depth maps with fewer errors.
@@
Robustness to Noise. Tab. 2 demonstrates that our method exhibits robustness in noisy environments. Similar to previous approaches [11, 47], we add Gaussian noise (mean 00 and standard deviation $0.07$) and Gaussian blur (standard deviation $3.6$) to the LR depth as new input. We can see that DORNet outperforms SFG [47] by $0.40cm$ in RMSE on the RGB-D-D dataset and surpasses SUFT by $0.61cm$ in RMSE on the TOFDSR [24] dataset.

### v4

@@ 1 Introduction
Blind depth super-resolution (DSR) aims to recover precise high-resolution (HR) depth from low-resolution (LR) depth with unknown degradation, which has been widely applied in many fields, such as virtual reality [17, 44, 33, 35], augmented reality [41, 31, 6, 48, 43], and 3D reconstruction [34, 3, 4, 46, 49]. As shown in Fig. 1(a), recent RGB-guided DSR methods [52, 58, 45, 26, 2] have been proposed that integrate RGB features aligned with input depth based on the assumption of known and fixed degradation, achieving excellent performance on synthetic data.
However, due to limitations in sensor technology and imaging environments, depth data captured from real-world scenes often suffer from unconventional and unknown degradation [47] (e.g., structural distortion and blur). Such real-world degradation results in structure inconsistency between depth and RGB, impairing the performance of previous methods. Moreover, real-world degradation labels are unavailable, preventing us from explicitly estimating the degradation between LR depth and HR depth.
@@
To address these issues, as shown in Fig. 1(b), we propose a degradation oriented and regularized network (DORNet). The DORNet utilizes degradation representations to guide the restoration of HR depth from real-world scenarios with unknown degradation. To this end, we present a self-supervised degradation learning strategy to estimate the implicit degradation representations between LR and HR depth. In this strategy, a router mechanism is first introduced to dynamically control the generation of degradation kernels with varying scales. We then design degradation regularization that leverages these kernels to deteriorate the predicted HR depth, yielding a new degraded depth. Consequently, the entire degradation process is learned by narrowing the distance between the new degraded depth and the LR depth, without using degradation labels. Furthermore, we observe that RGB can provide sharp and complete details for the degradation areas of the LR depth. Therefore, we propose utilizing the estimated degradation to adaptively select RGB features to guide and facilitate the RGB-D fusion. Concretely, we develop a degradation-oriented fusion scheme, deploying a degradation-oriented feature transformation module (DOFT). The DOFT produces filter kernels from learned degradation and then filters the RGB features, offering complementary contents for the depth features.
@@
Figure 3: Overview of DORNet. Given $\boldsymbol{D}_{up}$ as input, the degradation learning first encodes it to produce degradation representations $\boldsymbol{\tilde{D}}$ and $\boldsymbol{D}$. Then, $\boldsymbol{\tilde{D}}$, $\boldsymbol{D}$, $\boldsymbol{D}_{lr}$, and $\boldsymbol{I}$ are fed into multiple degradation-oriented feature transformation (DOFT) modules, generating the HR depth $\boldsymbol{D}_{hr}$. Finally, $\boldsymbol{D}$ and $\boldsymbol{D}_{hr}$ are sent to the degradation regularization to obtain $\boldsymbol{D}_{d}$, which is used as input for the degradation loss $\mathcal{L}_{deg}$ and the contrastive loss $\mathcal{L}_{cont}$. The degradation regularization only applies during training and adds no extra overhead in testing.
@@ 3.1 Network Architecture
Given LR depth $\boldsymbol{D}_{lr}\in R^{h\times w\times 1}$ with unknown degradation and RGB $\boldsymbol{I}\in R^{sh\times sw\times 3}$ as inputs, our method aims to recover accurate HR depth $\boldsymbol{D}_{hr}\in R^{sh\times sw\times 1}$ by learning the degradation representations. $h$, $w$, and $s$ represent the height, width, and upsampling factor, respectively.
As shown in Fig. 3, our DORNet mainly consists of a self-supervised degradation learning strategy (green part) and a degradation-oriented fusion scheme (orange part). Specifically, the upsampled LR depth $\boldsymbol{D}_{up}\in R^{sh\times sw\times 1}$ is first input into the degradation learning, producing both the router and the degradation representations, $\boldsymbol{\tilde{D}}$ and $\boldsymbol{D}$. Then, $\boldsymbol{\tilde{D}}$, $\boldsymbol{D}$, $\boldsymbol{D}_{lr}$, and $\boldsymbol{I}$ are sent to multiple degradation-oriented feature transformation modules (DOFT), which selectively propagate RGB information into the depth features, resulting in HR depth $\boldsymbol{D}_{hr}$. Next, the degradation regularization takes $\boldsymbol{D}$ as input and utilizes routing selection to adaptively generate degradation kernels with varying scales, all of which are sent into the filtering and summation modules together with $\boldsymbol{D}_{hr}$, obtaining the new degraded depth $\boldsymbol{D}_{d}$. Finally, $\boldsymbol{D}_{d}$ is employed as input for the degradation loss $\mathcal{L}_{deg}$ and the contrastive loss [39] $\mathcal{L}_{cont}$, further promoting the learning of degradation representations.
@@ 3.3 Degradation-Oriented Fusion
As shown in the orange part of Fig. 3, $\boldsymbol{D}_{lr}$ is first input into bicubic upsampling. Then, the upsampled LR depth and $\boldsymbol{I}$ are mapped to $\boldsymbol{F}_{d}^{0}$ and $\boldsymbol{F}_{r}^{0}$, respectively.
@@ 4.1 Experimental Setups
Datasets. We conduct extensive experiments on both real-world RGB-D-D [9], TOFDSR [46], and synthetic NYU-v2 [27] datasets. Specifically, for the RGB-D-D, the training set comprises $2,215$ RGB-D pairs, while the test set contains $405$ pairs. Additionally, the colorization method [14] is used to fill in the raw LR depth of the TOFDC [46], obtaining the TOFDSR that includes $10K$ RGB-D pairs for training and $560$ pairs for testing. In the real-world scenarios, the LR depth is obtained using the ToF camera of the Huawei P30 Pro. Following [12, 56, 37], the synthetic NYU-v2 consists of $1,000$ RGB-D pairs for training and $449$ pairs for testing, with the LR depth generated by bicubic downsampling from the GT depth.
@@ 4.2 Comparison with the State-of-the-Art
Furthermore, Figs. 8 and 9 present the visual results on the RGB-D-D and TOFDSR. In the error maps, a brighter color means a larger error. Obviously, for severely degraded LR depth, our method succeeds in recovering accurate depth structures. For instance, the handbag in Fig. 8 predicted by our method is more precise than others. Additionally, the error maps in Fig. 9 show that DORNet reconstructs HR depth with fewer errors.
@@
Robustness to Noise. Tab. 2 demonstrates that our method exhibits robustness in noisy environments. Similar to previous approaches [12, 53], we add Gaussian noise (mean 00 and standard deviation $0.07$) and Gaussian blur (standard deviation $3.6$) to upsampled LR depth as new input. We can see that DORNet outperforms SFG [53] by $0.40cm$ in RMSE on the RGB-D-D. For experiments on adding noise before LR depth pre-upsampling, please see our appendix.
@@ 5 Conclusion
Acknowledgements
This work was supported by the National Science Fund of China under Grant Nos. U24A20330 and 62361166670.